OpenStudy (anonymous):

Help please 2. Identify the equation(s) that are wrong. Explain any incorrect ones; provide complete details. A. log 6(m + n) = log 6 • log (m + n) B. log 10 = 1 ( I know this is correct) C. log m/log n = log(m/n)

OpenStudy (mathstudent55):

Apply this rule to the left side of A. What do you get? \(\log ab = \log a + \log b\)

OpenStudy (mathstudent55):

B. is correct as you stated.

OpenStudy (mathstudent55):

C. \(\dfrac{\log a}{\log b} = \log a - \log b\) Apply this rule to the left side of C.

OpenStudy (anonymous):

A) log 6 m * log 6 n C) log m - log n Are these correct?

OpenStudy (mathstudent55):

For A., the rule is the log of a product is the sum of the logs. The product is 6 * (m + n), so \(\log 6(m + n) = \log 6 + \log (m + n)\) This is not the equation of problem A., so A. is incorrect.

OpenStudy (mathstudent55):

If in A., you decide to multiply out (distribute) the 6, then you have: \(\log (6m + 6n)\) Now you have the log of a sum (the sum of 6m and 6n). There is no rule for the log of a sum.

OpenStudy (mathstudent55):

In C. you are correct. The equation should have been \(\log \dfrac{m}{n} = \log m - \log n\).

OpenStudy (anonymous):

Thank you (A) was the most confusing for me but I think I understand what you explained. It would just be log 6 (m +n) = log 6 (m) + log 6(n)

OpenStudy (mathstudent55):

No. That is still incorrect.

OpenStudy (mathstudent55):

Remember that the rule of simplifying the log of a product only works if you take the log of a product. Here are some simple examples: \(\log ab = \log a + \log b\) \(\log 6x = \log 6 + \log x\) \(log (5 \times 10) = \log 5 + \log 10\)

OpenStudy (mathstudent55):

Here is a slightly more complicated example: \(\large \log [(x + 1)(x - 5)] = \log (x + 1) + \log (x - 5) \)

OpenStudy (mathstudent55):

Notice that in all cases above, the log of the product was turned into the sum of the logs of the factors of the product.

OpenStudy (mathstudent55):

Now let's go back to A. in your problem above.

OpenStudy (mathstudent55):

This is what you were asked about. The question is whether it is a correct application of the rule of the log of a product or not. A. log 6(m + n) = log 6 • log (m + n)

OpenStudy (mathstudent55):

To find out, we apply the rule to the left side. The left side is: \(\log 6(m+ n)\) We need to take the log of the product of 6 and (m + n)

OpenStudy (mathstudent55):

Since the rule is that the log of a product is the sum of the logs, then that means that \(\log 6(m + n) = \log 6 + \log (m + n)\) Notice on the right side, we have the sum of log 6 and log (m + n) The problem is incorrect because is shows the product of log 6 and log (m + n)

OpenStudy (anonymous):

Oh! ok I see what you mean thank you.

OpenStudy (mathstudent55):

You're welcome.